OpenStudy (anonymous):

c. has an x-intercept of 6 and a y-intercept of 3 Equation: _____________________________ Explain why you chose the form you used. ________________________________________

OpenStudy (anonymous):

two points are (6,0) and (0,3) slope is (3-0)/(0-6) = -0.5 y = mx +c m is the slope c is the y intercept y = -0.5 +3